Jaybtw:

Write an expression for the area of square 5 by combining the areas of the four triangles and the one square.

Sherixn:

Ok so triangle = 1/2(base*height)

Sherixn:

We will use a and b to represent base and height

Sherixn:

We will multiply this by 4 because there are 4 triangles: 4*1/2(a*b) = 2ab

Jaybtw:

like how u teach me

Sherixn:

use the Pythagorean theorem a^2 + b^2 + c^2

Jaybtw:

thank you so much

Jaybtw:

u have help me out alot

Sherixn:

a^2 + 2ab+ b^2 = 2ab + c^2

Jaybtw:

so that is the answer

Sherixn:

The answer is 2ab + c^2

Sherixn:

@jaybtw wrote:
thank you so much
No problem (: